OpenStudy (anonymous):

Estimate the area of a square floor if each side is 12 ft 9 in long. Use whole numbers

OpenStudy (anonymous):

12 * 12 = 144 and 9 * 9 = 81. 81 / 12 = 6 with a remainder of 9. 144 + 6 = 150. Then just add the remaining 9 inches. 150 ft and 9 in.

OpenStudy (anonymous):

Estimate: A = 13² = 169 ft²